OpenStudy (anonymous):

Which of the following is a factor of the polynomial below? A. x – 2 B. x + 2 C. x + 5 D. x + 8

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

8x^2+24x+16

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

no thats the polynomial for the question

OpenStudy (cwrw238):

8x^2+24x+16 = 4(2x^2 + 6x + 4) = 4(2x +2)(x + 2)

OpenStudy (campbell_st):

an easy method is to use the factor theorem if x -2 is a factor then x = 2 is a zero so find P(2) BY substituting x = 2 into the polynomial if P(2) = 0.... then x - 2 is a factor repeat the process for the other factors... and it saves trying to factor or solve the quadratic...

OpenStudy (campbell_st):

so \[P(2) = 8(2)^2 + 24(2) + 16\]